# Martian time-series unraveled: A multi-scale nested approach with factorial variational autoencoders

### v4

@@ Abstract
Index Terms:
unsupervised learning, clustering, source separation, multi-scale, variational autoencoders.
@@ I Introduction
Figure 1: A schematic diagram depicting the proposed unsupervised multi-scale clustering and source separation framework, consisting of three stages. In the initial stage, indicated on the left in Fig. 1, we convert the data into the pyramidal scattering spectra representation space. The objective of this stage is to represent the data across significantly diverse timescales, ranging from less than a minute to an hour. The second stage involves training a generative model, namely a factorial variational encoder, which learns a joint probabilistic description of the multi-scale representations (indicated in the middle of Fig. 1 with arrows going from one block to another indicating representations associated with different timescales). The goal of this stage is to identify prominent sources in the dataset across several timescales, which will, in turn, be used as “prior information” in source separation. In the final stage, indicated on the right in Fig. 1, we separate the prominent source signals identified by the fVAE in the previous stage via solving an optimization problem in the scattering spectra space with the goal of separating a specific source of interest in the time domain. The code for partially reproducing the results can be found on GitHub.
@@
To demonstrate the applicability of our approach, we apply our approach to the full seismic dataset recorded by a seismometer on Mars during NASA’s Interior Exploration using Seismic Investigations, Geodesy and Heat Transport (InSight) mission [41, 42, 43]. The InSight lander’s seismometer—known as the SEIS instrument—detected marsquakes [44, 45, 46, 47] and transient atmospheric signals, such as wind and temperature changes, that provide information about the Martian atmosphere [48] and enable studying the interior structure and composition of the Red Planet [49]. The signal recorded by the InSight seismometer is heavily influenced by atmospheric activity and surface temperature [50, 51]. These non-seismic signals largely vary in timescale from minutes to hours and it would be crucial to capture this in multiple timescale. In addition to providing examples on the seismic dataset recorded during the InSight mission, we demonstrate the results of our approach in a controlled setting when applied to a stylized example involving multi-scale clustering and source separation of time series consisting of sources with three different timescale (cf. appendix D of the Appendix).
@@
Through analysis of this complex dataset via our approach, we observe that clusters identified at the finest timescale reveal distinct features such as glitches with and without precursors, with the former showing increased occurrence around Martian sunset. As the timescale broadens, clusters exhibit longer-duration events like oscillatory signals and bursts of high-frequency energy, possibly associated with wind dynamics. Moreover, coarser timescales unveil phenomena like wind gusts and atmospheric interactions, evident from characteristic waveforms and their occurrence patterns aligned with Martian day-night cycles. By leveraging clusters of glitches and wind-burst noises, we applied our source separation method and successfully separated these sources from the given time series. The results, illustrated through waveform visualizations, indicate successful separation with minimal alterations to the original waveform, validating the effectiveness of the approach. Additionally, as we show in appendix C of the Appendix, our method is also capable of separating background noise and glitches from a marsquake, which indicates the applicability of our method in cleaning sources that are not prominent enough to have their own clusters. Finally, the exploration of latent space clusters via the fVAE architecture provides insights into the learned representation of the dataset. Through low-dimensional visualizations, we examined clusters across different timescales. We observed that clusters in finer timescales exhibit clearer boundaries, suggesting more structured data. Moreover, the distribution of quality “A” broadband events and pressure drops across different timescales highlights the multi-scale representation of data captured by the fVAE architecture. Broadband events concentrate more in coarser timescales, while pressure drops are compactly situated in finer timescales, indicating the efficacy of the learned representation in capturing diverse temporal phenomena. Overall, these analyses offer valuable insights into the separation of prominent sources and the structure of latent space clusters, showcasing the capabilities of the proposed methods in analyzing complex datasets.
@@ II-A Wavelet scattering networks
[…]
@@ II-B Capturing non-Gaussian properties through scale dependencies
Sources $\mathbf{x}$ encountered in time-series studied in this paper can be considered are very often non-Gaussian processes. For sake of simplicity, let us assume $\mathbf{x}$ is a stationary source. If $\mathbf{x}$ were Gaussian then the different scale channels of a scattering network would be independent, however it is not the case in practice and these dependencies were shown to be crucial to characterize the non-Gaussian stochastic structure of $\mathbf{x}$ [35]. Such dependencies can be captured by considering the correlation matrix ${\mathbb{E}}[S(\mathbf{x})S(\mathbf{x})^{\top}]$:
@@ II-C Pyramidal averaging
Non-stationarity in the data $\mathbf{x}$ is in part explained by the presence of sources at different timescales. Our representation $\Psi(\mathbf{x})$ averages scattering spectra features on a certain window. If a source has a time-duration that is much smaller than the average window size, it will be averaged out, and the representation will contain little information about such a source. To take into account the variety of source timescales, we consider different averaging sizes in a causal manner. We replace $\underset{}{\operatorname{Ave}}\,{}$ in 2 by a multi-scale average pooling operator $\underset{}{\operatorname{Ave}}\,=\big{(}\underset{t\in w_{1}}{\operatorname{Ave}}\,,\ldots,\underset{t\in w_{K}}{\operatorname{Ave}}\,\big{)}$. The windows $w_{1}\subset\ldots\subset w_{K}$ have a pyramidal structure, they are of increasing size, and all ending at the same time, where $\underset{t\in w_{1}}{\operatorname{Ave}}\,$ considers recent past while $\underset{t\in w_{K}}{\operatorname{Ave}}\,$ considers distant past. To cover a large range of timescales, we choose $w_{k+1}$ to be four times longer than $w_{k}$. This defines a pyramidal scattering spectra representation $\Psi(x)=(\Psi_{1}(x),\ldots,\Psi_{K}(x))$ whose factor $k$ is $\Psi_{k}(\mathbf{x}):=\underset{t\in w_{k}}{\operatorname{Ave}}\,\big{(}S(\mathbf{x}),\operatorname{diag}\left(S(\mathbf{x})S(\mathbf{x})^{\top}\right)\big{)}$. The representation $\Psi(\mathbf{x})$ decomposes the variation in the stochastic structure of process $\mathbf{x}$ over time, through a multi-scale operator $\underset{}{\operatorname{Ave}}\,$.
@@ III-A Generative model
$p_{\theta}(\mathbf{u},\mathbf{y},\mathbf{z})$ $=p_{\theta}(\mathbf{u}\mid\mathbf{z})p_{\theta}(\mathbf{z}\mid\mathbf{y})p_{\theta}(\mathbf{y})$ (3a)
$=\prod_{i=0}^{s-1}p_{\theta}(\mathbf{u}_{i}\mid\mathbf{z}_{i})\,p_{\theta}(\mathbf{z}_{i}\mid y_{i})\,p_{\theta}(y_{i}).$ (3b)
@@
$p_{\theta}(\mathbf{z}_{i}\mid y_{i})$ $=\mathcal{N}\left(\mathbf{z}_{i}\mid\boldsymbol{\mu}_{z,i}(y;\boldsymbol{\theta}),\operatorname{diag}(\boldsymbol{\sigma}_{z,i}^{2}(y_{i};\boldsymbol{\theta}))\right),$ (4b)
$p_{\theta}(\mathbf{u}_{i}\mid\mathbf{z}_{i})$ $=\mathcal{N}\left(\mathbf{u}_{i}\mid\boldsymbol{\mu}_{u,i}(\mathbf{z}_{i};\boldsymbol{\theta}),\operatorname{diag}(\boldsymbol{\sigma}_{u,i}^{2}(\mathbf{z}_{i};\boldsymbol{\theta}))\right),$ (4c)
where $c_{i}$ represents the number of components in the Gaussian mixture latent distribution. $p_{\theta}(\mathbf{z}_{i}\mid y_{i})$ is modeled as a Gaussian distribution with $\boldsymbol{\mu}_{z,i}(y;\boldsymbol{\theta})$ and $\boldsymbol{\sigma}_{z,i}(y_{i};\boldsymbol{\theta})$ simply being learnable vectors for each $y_{i}\in\{0,\ldots,c_{i}-1\}$, which amounts to a Gaussian mixture model for $\mathbf{z}_{i}$. Conditioned on $\mathbf{z}_{i}$, $\mathbf{u}_{i}$ is also modeled as a Gaussian distribution with mean $\boldsymbol{\mu}_{u,i}(\mathbf{z}_{i};\boldsymbol{\theta})$ and diagonal covariance $\boldsymbol{\sigma}_{u,i}(\mathbf{z}_{i};\boldsymbol{\theta})$ parameterized using deep nets. The generative model setup outlined above translates to having independent decoders—i.e., mappings from latent variables to scattering spectra representations—for each timescale. This approach enables the independent synthesis of scattering spectra representations for each timescale for the downstream source separation step. To train this generative model tractably using VI, we define an inference model in the next section that approximates the latent posterior distribution.
@@ III-B Inference model
Evaluating the likelihood of the parametric distribution $p_{\theta}(\mathbf{u})$ requires marginalizing out the Gaussian mixture and the categorical latent variables from the joint distribution in 3. Unfortunately, this process is computationally infeasible due to the high-dimensionality of these distributions. To overcome this obstacle, we utilize amortized VI to approximate the latent posterior distribution $q(\mathbf{y},\mathbf{z}\mid\mathbf{u})$. This approximation uses the Evidence Lower Bound (ELBO) [37, 58] to approximate the model likelihood conditioned on the latent posterior on the wavelet scattering spectra representations for all timescales. To enable multi-scale clustering of the input data, we employ the following factorization of the posterior distribution,
$q_{\phi}(\mathbf{z},\mathbf{y}\mid\mathbf{u})$ $=q_{\phi}(\mathbf{z}\mid\mathbf{y},\mathbf{u})q_{\phi}(\mathbf{y}\mid\mathbf{u})$ (5a)
$=\prod_{i=1}^{s}q_{\phi}(\mathbf{z}_{i}\mid y_{i},\mathbf{u})q_{\phi}(y_{i}\mid\mathbf{u}).$ (5b)
@@
$q_{\phi}(y_{i}\mid\mathbf{u})$ $=\text{Cat}\left(\boldsymbol{\pi}_{i}(\mathbf{u};\boldsymbol{\phi})\right),$ (6a)
$q_{\phi}(\mathbf{z}_{i}\mid y_{i},\mathbf{u})$ $=\mathcal{N}\big{(}\mathbf{z}_{i}\mid\boldsymbol{\mu}_{z,i}(\mathbf{u},y_{i};\boldsymbol{\phi}),$ (6b)
$\quad\quad\quad\quad\ \operatorname{diag}(\boldsymbol{\sigma}_{z,i}^{2}(\mathbf{u},y_{i};\boldsymbol{\phi}))\big{)},$ (6c)
@@ III-C Training objective function
$\mathbb{KL}\,\big{(}p(\mathbf{u})\mid\mid p_{\theta}(\mathbf{u})\big{)}$ $=\operatorname*{\mathbb{E}}_{\mathbf{u}\sim p(\mathbf{u})}\big{[}\underbrace{\log p(\mathbf{u})}_{\text{const. w.r.t. }\boldsymbol{\theta}}-\log p_{\theta}(\mathbf{u})\big{]}$ (7a)
$=\operatorname*{\mathbb{E}}_{\mathbf{u}\sim p(\mathbf{u})}.$ (7b)
Evaluating the likelihood $p_{\theta}(\mathbf{u})$ is intractable due to the required marginalization over $\mathbf{y}$ and $\mathbf{z}$ (cf. 3). As a result, we approximate the likelihood with ELBO, which amounts to computing the expectation of joint distribution in 3 with respect to the latent posterior distribution, leading to the following training optimization problem:
@@
$\leq$ $\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\,\operatorname*{\mathbb{E}}_{\mathbf{u}\sim p(\mathbf{u})}\,\operatorname*{\mathbb{E}}_{\mathbf{z},\mathbf{y}\sim q_{\phi}(\mathbf{z},\mathbf{y}\mid\mathbf{u})}\Big{[}-\log\frac{p_{\theta}(\mathbf{u},\mathbf{y},\mathbf{z})}{q_{\phi}(\mathbf{z},\mathbf{y}\mid\mathbf{u})}\Big{]}$ (8b)
$=$ $\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\,\sum_{i=1}^{s}\operatorname*{\mathbb{E}}_{\mathbf{u}\sim p(\mathbf{u})}\,\Big{[}\operatorname*{\mathbb{E}}_{\mathbf{z}_{i}\sim q_{\phi}(\mathbf{z}_{i}\mid y_{i},\mathbf{u})}\,\underbrace{\big{[}-\log p_{\theta}(\mathbf{u}_{i}\mid\mathbf{z}_{i})\big{]}}_{\begin{subarray}{c}\text{Per-scale reconstruction loss}\end{subarray}}$ (8c)
$+\underbrace{\mathbb{KL}\,\big{(}q_{\phi}(y_{i}\mid\mathbf{u})\mid\mid p_{\theta}(y_{i})\big{)}}_{\begin{subarray}{c}\text{Categorical prioron }y_{i}\end{subarray}}$ (8d)
$+\operatorname*{\mathbb{E}}_{y_{i}\sim q_{\phi}(y_{i}\mid\mathbf{u})}\,\underbrace{\big{[}\mathbb{KL}\,\big{(}q_{\phi}(\mathbf{z}_{i}\mid y_{i},\mathbf{u})\mid\mid p_{\theta}(\mathbf{z}_{i}\mid y_{i})\big{)}\big{]}}_{\begin{subarray}{c}\text{Gaussian mixture prior on }\mathbf{z}_{i}\end{subarray}}\Big{]}.$ (8e)
The expectations in the optimization problem above can be approximated using Monte Carlo integration over samples from their respective distributions.

### v5

# Multi-scale clustering and source separation of InSight mission seismic data

@@ I Introduction
Figure 1: A schematic diagram depicting the proposed unsupervised multi-scale clustering and source separation framework, consisting of three stages. In the initial stage, indicated on the left in Fig. 1, we convert the data into the pyramidal scattering spectra representation space. The objective of this stage is to represent the data across significantly diverse timescales, ranging from less than a minute to an hour. The second stage involves training a generative model, namely a factorial variational encoder, which learns a joint probabilistic description of the multi-scale representations (indicated in the middle of Fig. 1 with arrows going from one block to another indicating representations associated with different timescales). The goal of this stage is to identify prominent sources in the dataset across several timescales, which will, in turn, be used as “prior information” in source separation. In the final stage, indicated on the right in Fig. 1, we separate the prominent source signals identified by the fVAE in the previous stage via solving an optimization problem in the scattering spectra space with the goal of separating a specific source of interest in the time domain. appendix A of the Appendix provides a more detailed walkthrough of the proposed framework. The code for partially reproducing the results will be made available after the review process.
@@
To demonstrate the applicability of our approach, we apply our approach to the entire seismic dataset recorded by a seismometer on Mars during NASA’s Interior Exploration using Seismic Investigations, Geodesy and Heat Transport (InSight) mission [41, 42, 43]. The InSight lander’s seismometer—known as the SEIS instrument—detected marsquakes [44, 45, 46, 47] and transient atmospheric signals, such as wind and temperature changes, that provide information about the Martian atmosphere [48] and enable studying the interior structure and composition of the Red Planet [49]. The signal recorded by the InSight seismometer is heavily influenced by atmospheric activity and surface temperature [50, 51]. These non-seismic signals largely vary in timescale from minutes to hours and it would be crucial to capture this in multiple timescale. In addition to providing examples on the seismic dataset recorded during the InSight mission, we demonstrate the results of our approach in a controlled setting when applied to a stylized example involving multi-scale clustering and source separation of time series consisting of sources with three different timescale (cf. appendix I of the Appendix).
@@
Through analysis of this complex dataset via our approach, we observe that clusters identified at the finest timescale reveal distinct features such as glitches with and without precursors, with the former showing increased occurrence around Martian sunset. As the timescale broadens, clusters exhibit longer-duration events like oscillatory signals and bursts of high-frequency energy, possibly associated with wind dynamics. Moreover, coarser timescales unveil phenomena like wind gusts and atmospheric interactions, evident from characteristic waveforms and their occurrence patterns aligned with Martian day-night cycles. By leveraging clusters of glitches and wind-burst noises, we applied our source separation method and successfully separated these sources from the given time series. The results, illustrated through waveform visualizations, indicate successful separation with minimal alterations to the original waveform, validating the effectiveness of the approach. Additionally, as we show in appendix H of the Appendix, our method is also capable of separating background noise and glitches from a marsquake, which indicates the applicability of our method in cleaning sources that are not prominent enough to have their own clusters. Finally, the exploration of latent space clusters via the fVAE architecture provides insights into the learned representation of the dataset. Through low-dimensional visualizations, we examined clusters across different timescales. We observed that clusters in finer timescales exhibit clearer boundaries, suggesting more structured data. Moreover, the distribution of quality “A” broadband events and pressure drops across different timescales highlights the multi-scale representation of data captured by the fVAE architecture. Broadband events concentrate more in coarser timescales, while pressure drops are compactly situated in finer timescales, indicating the efficacy of the learned representation in capturing diverse temporal phenomena. Overall, these analyses offer valuable insights into the separation of prominent sources and the structure of latent space clusters, showcasing the capabilities of the proposed methods in analyzing complex datasets.
@@ II-A Wavelet scattering networks
A scattering network [36] is a cascade of wavelet operators $\mathbf{W}$ followed by nonlinear activation functions (akin to a typical convolutional neural network). A wavelet transform operator $\mathbf{W}$ is a convolutional operator with predefined wavelet filters that extracts variations at separate scales. These filters include a low-pass filter $\varphi_{J}(t)$ and $J$ complex-valued band-pass filters $\psi_{j}(t)=2^{-j}\psi(2^{-j}t),\ 1\leq j\leq J$, which are obtained by the dilation of a mother wavelet $\psi(t)$ that have zero mean and a fast decay away from $t=0$. The wavelet coefficient $\mathbf{W}\mathbf{x}(t,j)=x\star\psi_{j}(t)$ extracts variations of the input signal $\mathbf{x}(t)$ around time $t$ at scale $2^{j}$. In order to characterize time-evolution of the wavelet coefficients e.g., envelope modulation, we apply a modulus $|\cdot|$ and cascade a second wavelet operator. The output of a two-layer scattering network $S$ is $S(\mathbf{x}):=\big(\mathbf{W}\mathbf{x}\,,\,\mathbf{W}|\mathbf{W}\mathbf{x}|\big)^{\top}$, it extracts variations of signal $\mathbf{x}$ and its multi-scale envelopes $|\mathbf{W}\mathbf{x}|$ at different times and different scales. Even though these networks have been successfully employed in tasks such as intermittency analysis [52], clustering [53], event detection and segmentation [54] (with learnable wavelets), they are not sufficient to build an accurate description of a multi-scale process, as they fail to capture crucial dependencies across different scales [35].
The architecture we use for wavelet scattering spectra computation throughout this paper is a two-layer scattering network with $J=8$ different octaves with one wavelet per octave.
@@ II-B Capturing non-Gaussian properties through scale dependencies
Sources $\mathbf{x}$ encountered in time-series studied in this paper are very often non-Gaussian processes. For sake of simplicity, let us assume $\mathbf{x}$ is a stationary source. If $\mathbf{x}$ were Gaussian then the different scale channels of a scattering network would be independent, however it is not the case in practice and these dependencies were shown to be crucial to characterize the non-Gaussian stochastic structure of $\mathbf{x}$ [35]. Such dependencies can be captured by considering the correlation matrix ${\mathbb{E}}[S(\mathbf{x})S(\mathbf{x})^{\top}]$:
@@ II-C Pyramidal averaging
Non-stationarity in the data $\mathbf{x}$ is in part explained by the presence of sources at different timescales. Our representation $\Psi(\mathbf{x})$ averages scattering spectra features on a certain window. If a source has a time-duration that is much smaller than the average window size, it will be averaged out, and the representation will contain little information about such a source. To take into account the variety of source timescales, we consider different averaging sizes in a causal manner. We replace $\underset{}{\operatorname{Ave}}\,{}$ in 2 by a multi-scale average pooling operator $\underset{}{\operatorname{Ave}}\,=\big(\underset{t\in w_{1}}{\operatorname{Ave}}\,,\ldots,\underset{t\in w_{K}}{\operatorname{Ave}}\,\big)$. The windows $w_{1}\subset\ldots\subset w_{K}$ have a pyramidal structure, they are of increasing size, and all ending at the same time, where $\underset{t\in w_{1}}{\operatorname{Ave}}\,$ considers recent past while $\underset{t\in w_{K}}{\operatorname{Ave}}\,$ considers distant past. To cover a large range of timescales, we choose $w_{k+1}$ to be four times longer than $w_{k}$. This defines a pyramidal scattering spectra representation $\Psi(x)=(\Psi_{1}(x),\ldots,\Psi_{K}(x))$ whose factor $k$ is
$\Psi_{k}(\mathbf{x}):=\underset{t\in w_{k}}{\operatorname{Ave}}\,\big(S(\mathbf{x}),\operatorname{diag}\left(S(\mathbf{x})S(\mathbf{x})^{\top}\right)\big).$ (3)
The full representation $\Psi(\mathbf{x})$, consisting of all factors decomposes the variation in the stochastic structure of process $\mathbf{x}$ over time through the multi-scale pooling operator $\underset{}{\operatorname{Ave}}\,$.
@@ III-A Generative model
$p_{\theta}(\mathbf{u},\mathbf{y},\mathbf{z})$ $=p_{\theta}(\mathbf{u}|\mathbf{z})p_{\theta}(\mathbf{z}|\mathbf{y})p_{\theta}(\mathbf{y})$ (4a)
$=\prod_{i=0}^{s-1}p_{\theta}(\mathbf{u}_{i}|\mathbf{z}_{i})\,p_{\theta}(\mathbf{z}_{i}|y_{i})\,p_{\theta}(y_{i}).$ (4b)
@@
$p_{\theta}(\mathbf{z}_{i}|y_{i})$ $=\mathcal{N}\left(\mathbf{z}_{i}|\boldsymbol{\mu}_{z,i}(y;\boldsymbol{\theta}),\operatorname{diag}(\boldsymbol{\sigma}_{z,i}^{2}(y_{i};\boldsymbol{\theta}))\right),$ (5b)
$p_{\theta}(\mathbf{u}_{i}|\mathbf{z}_{i})$ $=\mathcal{N}\left(\mathbf{u}_{i}|\boldsymbol{\mu}_{u,i}(\mathbf{z}_{i};\boldsymbol{\theta}),\operatorname{diag}(\boldsymbol{\sigma}_{u,i}^{2}(\mathbf{z}_{i};\boldsymbol{\theta}))\right),$ (5c)
where $c_{i}$ represents the number of components in the Gaussian mixture latent distribution. $p_{\theta}(\mathbf{z}_{i}|y_{i})$ is modeled as a Gaussian distribution with $\boldsymbol{\mu}_{z,i}(y;\boldsymbol{\theta})$ and $\boldsymbol{\sigma}_{z,i}(y_{i};\boldsymbol{\theta})$ simply being learnable vectors for each $y_{i}\in\{0,\ldots,c_{i}-1\}$, which amounts to a Gaussian mixture model for $\mathbf{z}_{i}$. Conditioned on $\mathbf{z}_{i}$, $\mathbf{u}_{i}$ is also modeled as a Gaussian distribution with mean $\boldsymbol{\mu}_{u,i}(\mathbf{z}_{i};\boldsymbol{\theta})$ and diagonal covariance $\boldsymbol{\sigma}_{u,i}(\mathbf{z}_{i};\boldsymbol{\theta})$ parameterized using deep nets. The generative model setup outlined above translates to having independent decoders—i.e., mappings from latent variables to scattering spectra representations—for each timescale. This approach enables the independent synthesis of scattering spectra representations for each timescale for the downstream source separation step. To train this generative model tractably using VI, we define an inference model in the next section that approximates the latent posterior distribution.
@@ III-B Inference model
Evaluating the likelihood of the parametric distribution $p_{\theta}(\mathbf{u})$ requires marginalizing out the Gaussian mixture and the categorical latent variables from the joint distribution in 4. Unfortunately, this process is computationally infeasible due to the high-dimensionality of these distributions. To overcome this obstacle, we utilize amortized VI to approximate the latent posterior distribution $q(\mathbf{y},\mathbf{z}|\mathbf{u})$. This approximation uses the Evidence Lower Bound (ELBO) [37, 58] to approximate the model likelihood conditioned on the latent posterior on the wavelet scattering spectra representations for all timescales. To enable multi-scale clustering of the input data, we employ the following factorization of the posterior distribution,
$q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})$ $=q_{\phi}(\mathbf{z}|\mathbf{y},\mathbf{u})q_{\phi}(\mathbf{y}|\mathbf{u})$ (6a)
$=\prod_{i=1}^{s}q_{\phi}(\mathbf{z}_{i}|y_{i},\mathbf{u})q_{\phi}(y_{i}|\mathbf{u}).$ (6b)
@@
$q_{\phi}(y_{i}|\mathbf{u})$ $=\text{Cat}\left(\boldsymbol{\pi}_{i}(\mathbf{u};\boldsymbol{\phi})\right),$ (7a)
$q_{\phi}(\mathbf{z}_{i}|y_{i},\mathbf{u})$ $=\mathcal{N}\big(\mathbf{z}_{i}|\boldsymbol{\mu}_{z,i}(\mathbf{u},y_{i};\boldsymbol{\phi}),$ (7b)
$\quad\quad\quad\ \ \operatorname{diag}(\boldsymbol{\sigma}_{z,i}^{2}(\mathbf{u},y_{i};\boldsymbol{\phi}))\big),$ (7c)
@@ III-C Training objective function
$\operatorname*{arg\,min}_{\boldsymbol{\theta}}\,\mathbb{KL}\,\big(p(\mathbf{u})\,||\,p_{\theta}(\mathbf{u})\big)$ (8a)
$=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\,\operatorname*{\mathbb{E}}_{\mathbf{u}\sim p(\mathbf{u})}\big[\underbrace{\log p(\mathbf{u})}_{\text{const. w.r.t. }\boldsymbol{\theta}}-\log p_{\theta}(\mathbf{u})\big]$ (8b)
$=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\,\operatorname*{\mathbb{E}}_{\mathbf{u}\sim p(\mathbf{u})}\big[-\log p_{\theta}(\mathbf{u})\big].$ (8c)
Evaluating the likelihood $p_{\theta}(\mathbf{u})$ is intractable due to the required marginalization over $\mathbf{y}$ and $\mathbf{z}$ (cf. 4). To overcome this intractability, we maximize a lower bound to the likelihood (or minimize the negative lower bound), commonly referred to as the ELBO. This lower bound to the likelihood is obtained by computing the expectation of joint distribution in 4 with respect to the latent posterior distribution, leading to the following training optimization problem:
@@
$\leq$ $\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\sum_{i=0}^{s-1}\operatorname*{\mathbb{E}}_{\mathbf{u}\sim p(\mathbf{u})}\,\Big[\operatorname*{\mathbb{E}}_{\begin{subarray}{c}y_{i}\sim q_{\phi}(y_{i}|\mathbf{u})\\[-1.0pt] \mathbf{z}_{i}\sim q_{\phi}(\mathbf{z}_{i}|y_{i},\mathbf{u})\end{subarray}}$ (9b)
$\quad\Big[-\log\frac{p_{\theta}(\mathbf{u}_{i}|\mathbf{z}_{i})\,p_{\theta}(\mathbf{z}_{i}|y_{i})\,p_{\theta}(y_{i})}{q_{\phi}(\mathbf{z}_{i}|y_{i},\mathbf{u})\,q_{\phi}(y_{i}|\mathbf{u})}\Big]\Big]$ (9c)
$=$ $\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\,\sum_{i=1}^{s}\Big[\operatorname*{\mathbb{E}}_{\begin{subarray}{c}\mathbf{u}\sim p(\mathbf{u})\\[-1.0pt] \mathbf{z}_{i}\sim q_{\phi}(\mathbf{z}_{i}|y_{i},\mathbf{u})\end{subarray}}\,\underbrace{\big[-\log p_{\theta}(\mathbf{u}_{i}|\mathbf{z}_{i})\big]}_{\begin{subarray}{c}\text{Per-scale reconstruction loss}\end{subarray}}$ (9d)
$+\operatorname*{\mathbb{E}}_{\mathbf{u}\sim p(\mathbf{u})}\,\underbrace{\mathbb{KL}\,\big(q_{\phi}(y_{i}|\mathbf{u})\,||\,p_{\theta}(y_{i})\big)}_{\begin{subarray}{c}\text{Categorical prior on }y_{i}\end{subarray}}$ (9e)
$+\operatorname*{\mathbb{E}}_{\begin{subarray}{c}\mathbf{u}\sim p(\mathbf{u})\\[-1.0pt] y_{i}\sim q_{\phi}(y_{i}|\mathbf{u})\end{subarray}}\,\underbrace{\big[\mathbb{KL}\,\big(q_{\phi}(\mathbf{z}_{i}|y_{i},\mathbf{u})\,||\,p_{\theta}(\mathbf{z}_{i}|y_{i})\big)\big]}_{\begin{subarray}{c}\text{Gaussian mixture prior on }\mathbf{z}_{i}\end{subarray}}\Big].$ (9f)
The expectations in the optimization problem above can be approximated using Monte Carlo integration over samples from their respective distributions. For a detailed derivation, please refer to appendix B of the Appendix.